(2.4192 + M)² = (1 + 0.2 m²) Please solue for M. Thank you.
Answer #1

Sol (2.41924M) (1 + 0.2M²) 2.4192 M = (1+0-214213. 5 2041924 = 13+162 m2 t 0.008 Mo + 0.6142 to. 1244 0.008 Mo + 0.12 44+ 12M